# L2T-Tune:LLM‑Guided Hybrid Database Tuning with LHS and TD3

### v1

@@ I Introduction
  • Stage Three: Random Forest (RF)[21]/Principal Components Analysis (PCA)[22] + TD3 fine-tuning. Using the Stage-One samples, RF selects impactful knobs and PCA compresses state features; we then perform full TD3 reinforcement learning fine-tuning to reach the final optimum.
During the reinforcement learning tuning process, 63 parameters are used to represent the current state of MySQL, which guides the actor’s actions in the TD3 model. These 63 parameters are reduced in dimensionality using Principal Component Analysis (PCA), further enhancing efficiency. Finally, the TD3 model is used for reinforcement learning training. The TD3 model continuously updates the actor and critic networks based on feedback from the environment, fine-tuning the database knobs for optimal performance.
@@
    Superior Tuning Performance: The experimental results show that compared with the most advanced methods, our approach achieves up to 73% performance improvement in TPC-C benchmarks, and it only requires 30 steps during the online tuning stage to reach the best effect.
@@ III-A Stage 1: LHS Warm-Start for Diverse Exploration
Beyond improving warm-start configurations, The stratified sampling of LHS ensures more uniform coverage across all knobs. This uniformity serves two key purposes: (1) it raises the performance baseline for subsequent DB-BERT (Stage 2) and TD3 (Stage 3), minimizing the impact of poor initializations, and (2) it creates a well-distributed dataset crucial for Random-Forest-based knob selection in Stage 3, allowing for robust feature-importance estimation without the bias toward clustered regions that GA tends to introduce. By combining enhanced warm-start performance with consistent data sampling for dimensionality reduction, LHS lays a strong foundation for L2T-Tune’s three-stage pipeline.
@@ III-B1 Improved GPTuner
Two-stage tuning. GPTuner then uses a coarse-to-fine search. In the coarse stage it builds a Tiny Feasible Space from documentation suggestions and range-scaled variants, seeds it with a small LHS design, fits a random-forest surrogate, and runs SMAC [36] (RF-based Bayesian optimization) for a few iterations. The fine stage then widens the search to the full heterogeneous ranges and continues SMAC, bootstrapped by the coarse samples, to refine numeric values. The pipeline has three parts: a Knowledge Handler that builds the Tuning Lake, a Search-Space Optimizer for knob selection and range shaping, and a Configuration Recommender that executes the coarse-to-fine loop.
@@ III-B2 DB-BERT
Online hint-guided tuning. In the online phase, DB-BERT treats each normalized hint for knob $k$ as a base value $v_{k}$ and lets an A2C controller refine it. At every step the policy selects, for each hinted knob, a multiplicative factor $f\in\{0.25,0.5,1,2,4\}$ to scale $v_{k}$ and an importance weight $w\in\{1,2,4,8,16\}$ to prioritize impactful hints; the final configuration is composed from the $f\cdot v_{k}$ values aggregated by $w$. The system applies the configuration, evaluates the workload, and uses the reward (e.g., $\mathrm{TPS}/\max(\mathrm{p95},1)$) to update the actor–critic. Constraining actions to $(f,w)$ over a small set of hinted knobs enables fast convergence in roughly a few dozen trials.
Given this discrete action design (five multipliers and five weights applied in alternating steps), we experimentally replaced A2C with DDPG and TD3 and observed no improvement. The result is that continuous actions must be snapped back to the same five levels, which nullifies the advantage of continuous control and makes the learning signal nearly piecewise constant. Therefore, we retain the original method of dbbert.
@@ III-C Stage 3: Dimensionality Reduction and TD3 Fine-Tuning
On top of the LHS warm start and the LLM-guided recommendations, we add a third stage for RL fine-tuning. Using the uniform samples collected in Stage 1 (the shared ((S,A,P)) pool), we compute knob importances with Random Forest and select the top-(K) knobs; we also fit PCA on the 63-dim state using the same Stage-1 traces to obtain a compact state representation. TD3 then operates on this reduced action/state space to refine the configuration.
III-C1 Dimensionality Reduction: Random Forest + PCA
@@
Random Forest knob selection.
MySQL 8.0 exposes $266$ tunable knobs (e.g., memory: innodb_buffer_pool_size, key_buffer_size; concurrency: max_connections, innodb_thread_concurrency; I/O: innodb_flush_log_at_trx_commit, sync_binlog). Empirically, returns diminish as the number of tuned knobs grows. We therefore perform Random Forest-based feature selection using the $120$ uniformly distributed LHS samples (Stage 1):
@@
    Random Forest regression: We use 100 trees to calculate feature importance, measuring the reduction in mean squared error (MSE) for each knob across the trees.
@@ IV Experiments
We first conducted experiments on the MySQL database under three different workloads: Sysbench read-only, Sysbench read-write (1:1), and Sysbench write-only, as well as a TPC-C read-write workload. Each experiment was performed with a 60-second stress test. The database setup consisted of 32 tables, each containing 100,000 rows. For Sysbench read-only, the test was run with 64 threads, while the other workloads used 32 threads. The evaluation metric for all experiments was $\mathrm{TPS}/\mathrm{p95}$. These experiments were designed to evaluate the performance of the tuning models under various database load conditions.
@@
Importantly, our GPTuner ablation shows that coarse-only (c) often reaches a strong configuration faster and sometimes higher than coarse+fine (c+f). The reason is that the coarse budget is too small, so fine optimization starts before coarse lands near a good basin, leading to premature, locally biased refinement. In L2T-Tune we deliberately use coarse as the Stage 2 recommender and delegate fine-grained adjustment to Stage 3 TD3; coupled with Stage 1 warm-start that filters knobs by hardware-aware priors, the RL agent operates in a compact, high-signal subspace and thus surpasses GPTuner in both speed and final quality.
Quantitatively (Table II), relative to the strongest non-RL baseline GPTuner (coarse+fine), our best L2T-Tune variant improves TPS/p95 by +22.6% on Sysbench Read (L2T-Tune(gp) 413.68 vs 377.34), +26.4% on Sysbench RW (L2T-Tune(db) 179.69 vs 142.19), +26.3% on Sysbench Write (L2T-Tune(db) 649.72 vs 514.56), and +73.0% on TPC-C RW (L2T-Tune(db) 58.45 vs 33.78), averaging a +37.1% uplift. Compared with HUNTER and CDBTune, inserting the LLM stage dramatically accelerates RL convergence; within the same budget on a single server, RL-only systems struggle to reach competitive levels.
@@
In response to this challenge, we have developed a novel semi-transfer approach based on our model. This method involves transferring only the hot-start component and subsequently leveraging dbbert to effect rapid improvement over a relatively short period. Specifically, in our experiment, we executed 15 steps.

### v2

@@ I Introduction
  • Stage Three: Random Forest (RF)[21]/Principal Components Analysis (PCA)[22] + TD3 fine-tuning. Using the Stage-One samples, RF selects impactful knobs and PCA compresses state features. We then perform full TD3 reinforcement learning fine-tuning to reach the final optimum.
During the reinforcement learning tuning process, 63 parameters are used to represent the current state of MySQL, which guides the actor’s actions in the TD3 model. These 63 parameters are reduced in dimensionality using PCA, further enhancing efficiency. Finally, the TD3 model is used for reinforcement learning training. The TD3 model continuously updates the actor and critic networks based on feedback from the environment, fine-tuning the database knobs for optimal performance.
@@
    Superior Tuning Performance: Across workloads, our approach achieves an average improvement of 37.1%, reaches up to 73% on TPC-C, and requires only about 30 online-tuning steps to reach the best configuration.
@@ III-A Stage 1: LHS Warm-Start for Diverse Exploration
Beyond improving warm-start configurations, the stratified sampling of LHS ensures more uniform coverage across all knobs. This uniformity serves two key purposes: (1) it raises the performance baseline for subsequent LLMs (Stage 2) and TD3 (Stage 3), minimizing the impact of poor initializations, and (2) it creates a well-distributed dataset crucial for Random-Forest-based knob selection in Stage 3, allowing for robust feature-importance estimation without the bias toward clustered regions that GA tends to introduce. By combining enhanced warm-start performance with consistent data sampling for dimensionality reduction, LHS lays a strong foundation for L2T-Tune’s three-stage pipeline.
@@ III-B1 Improved GPTuner
Two-stage tuning. GPTuner then uses a coarse-to-fine search. In the coarse stage, it builds a Tiny Feasible Space from documentation suggestions and range-scaled variants, seeds it with a small LHS design, fits a random-forest surrogate, and runs SMAC [36] (RF-based Bayesian optimization) for a few iterations. The fine stage then widens the search to the full heterogeneous ranges and continues SMAC, bootstrapped by the coarse samples, to refine numeric values. The pipeline has three parts: a Knowledge Handler that builds the Tuning Lake, a Search-Space Optimizer for knob selection and range shaping, and a Configuration Recommender that executes the coarse-to-fine loop.
@@ III-B2 DB-BERT
Online hint-guided tuning. In the online phase, DB-BERT treats each normalized hint for knob $k$ as a base value $v_{k}$ and lets an A2C controller refine it. At every step the policy selects, for each hinted knob, a multiplicative factor $f\in\{0.25,0.5,1,2,4\}$ to scale $v_{k}$ and an importance weight $w\in\{1,2,4,8,16\}$ to prioritize impactful hints; the final configuration is composed from the $f\cdot v_{k}$ values aggregated by $w$. The system applies the configuration, evaluates the workload, and uses the reward to update the actor–critic. Constraining actions to $(f,w)$ over a small set of hinted knobs enables fast convergence in roughly a few dozen trials.
Given this discrete action design (five multipliers and five weights applied in alternating steps), we experimentally replace A2C with DDPG and TD3 and observed no improvement. The result is that continuous actions must be snapped back to the same five levels, which nullifies the advantage of continuous control and makes the learning signal nearly piecewise constant. Therefore, we retain the original method of dbbert.
@@ III-C Stage 3: Dimensionality Reduction and TD3 Fine-Tuning
On top of the LHS warm start and the LLM-guided recommendations, we add a third stage for RL fine-tuning. Using the uniform samples collected in Stage 1 (the shared ((S,A,P)) pool), we compute knob importances with RF and select the top-(K) knobs; we also fit PCA on the 63-dim state using the same Stage-1 traces to obtain a compact state representation. TD3 then operates on this reduced action/state space to refine the configuration.
III-C1 Dimensionality Reduction: RF + PCA
@@
RF knob selection
MySQL 8.0 exposes $266$ tunable knobs (e.g., memory: innodb_buffer_pool_size, key_buffer_size; concurrency: max_connections, innodb_thread_concurrency; I/O: innodb_flush_log_at_trx_commit, sync_binlog). Empirically, returns diminish as the number of tuned knobs grows. We therefore perform RF-based feature selection using the $120$ uniformly distributed LHS samples (Stage 1):
@@
    RF regression: We use 100 trees to calculate feature importance, measuring the reduction in mean squared error (MSE) for each knob across the trees.
@@ IV Experiments
We first conducted experiments on the MySQL database under four different workloads: Sysbench read-only, Sysbench read-write (1:1), and Sysbench write-only, as well as a TPC-C read-write workload. Each experiment was performed with a 60-second stress test. The database setup consisted of 32 tables, each containing 100,000 rows. For Sysbench read-only, the test was run with 64 threads, while the other workloads used 32 threads. The evaluation metric for all experiments was $\mathrm{TPS}/\mathrm{p95}$. These experiments were designed to evaluate the performance of the tuning models under various database load conditions.
@@
Importantly, coarse-only (c) often reaches a strong configuration faster and sometimes higher than coarse+fine (c+f). The reason is that the coarse budget is too small; consequently, fine optimization starts before coarse lands near a good basin, leading to premature, locally biased refinement. In L2T-Tune we deliberately use coarse as the Stage 2 recommender and delegate fine-grained adjustment to Stage 3 TD3; coupled with Stage 1 warm-start that filters knobs by hardware-aware priors, the RL agent operates in a compact, high-signal subspace and thus surpasses GPTuner in both speed and final quality.
Quantitatively (Table II), relative to the strongest baseline GPTuner (coarse+fine), our best L2T-Tune variant improves TPS/p95 by +22.6% on Sysbench Read (L2T-Tune(gp) 413.68 vs 377.34), +26.4% on Sysbench RW (L2T-Tune(db) 179.69 vs 142.19), +26.3% on Sysbench Write (L2T-Tune(db) 649.72 vs 514.56), and +73.0% on TPC-C RW (L2T-Tune(db) 58.45 vs 33.78), averaging a +37.1% uplift. Compared with HUNTER and CDBTune, inserting the LLM stage dramatically accelerates RL convergence; within the same budget on a single server, RL-only systems struggle to reach competitive levels.
@@
In response to this challenge, we have developed a novel semi-transfer approach based on our model. This method involves transferring only the hot-start component and subsequently leveraging dbbert to effect rapid improvement over a relatively short period.